# X-Mobility: End-To-End Generalizable Navigation via World Modeling

### v2

@@ II-C2 Observation Encoding
To update the belief state, X-Mobility takes front-view image and robot states, including linear and angular velocities, as inputs. These inputs are embedded into a compact, low-dimensional representation to effectively learn the dynamics described in Eqn. (1). To capture visual features, we employ a pretrained DINOv2 [20] model as the image encoder. The image is embedded into a 1-D token $u_{t}\in\mathbb{R}^{768}$ by concatenating the class token with the average-pooled patch tokens. DINOv2’s strong feature representation and robust performance across various image conditions can enhance X-Mobility’s ability to generalize well across different environments in both simulation and real world. The linear and angular speeds are encoded into $m_{t}\in\mathbb{R}^{32}$ using fully connected layers. This speed embedding is then concatenated with the image embedding to produce the observation embedding $o_{t}=[u_{t},m_{t}]$ for state estimation.
@@ III-A Dataset
speed $\mathbb{R}^{2}$checkmark $\checkmark$
@@ V-A Sim2Real Transfer
Additionally, we also benchmarked the model’s run-time performance on the Jetson AGX Orin, as showing in Table V, underscoring X-Mobility’s high computational efficiency for edge-device navigation.
TABLE V: Inference Latency and GPU Memory Usage
@@ Appendix A Model Description
The model parameters of X-Mobility are listed in Table VI, while Table VII outlines the hyperparameters used for training.
TABLE VI: Model parameters

### v3

@@ II-C2 Observation Encoding
To update the belief state, X-Mobility takes front-view image and robot states, including linear forward velocity, as inputs. These inputs are embedded into a compact, low-dimensional representation to effectively learn the dynamics described in Eqn. (1). To capture visual features, we employ a pretrained DINOv2 [20] model as the image encoder. The image is embedded into a 1-D token $u_{t}\in\mathbb{R}^{768}$ by concatenating the class token with the average-pooled patch tokens. DINOv2’s strong feature representation and robust performance across various image conditions can enhance X-Mobility’s ability to generalize well across different environments in both simulation and real world. The linear forward speed is encoded into $m_{t}\in\mathbb{R}^{32}$ using fully connected layers. This speed embedding is then concatenated with the image embedding to produce the observation embedding $o_{t}=[u_{t},m_{t}]$ for state estimation.
@@ III-A Dataset
speed $\mathbb{R}^{1}$checkmark $\checkmark$
@@ V-A Sim2Real Transfer
To more rigorously evaluate real-world performance, we set up environments with varying obstacle configurations and lighting conditions (see Fig. 7). The final results, summarized in Table V, demonstrate performance comparable to the benchmarks observed in simulation. Additionally, we also benchmarked the model’s run-time performance on the Jetson AGX Orin, as showing in Table VI, underscoring X-Mobility’s high computational efficiency for edge-device navigation.
Figure 7: Environments setup for benchmark on real robot.
TABLE V: Benchmark on real robot
Scene Lightning SR AA($rad/s^{2}$)
Single Obstacle Normal 10/10 0.267
Single Obstacle Dark 10/10 0.267
Multi Obstacles Normal 8/9 0.205
Multi Obstacles Dark 8/9 0.273
TABLE VI: Inference Latency and GPU Memory Usage
@@ Appendix A Model Description
The model parameters of X-Mobility are listed in Table VII, while Table VIII outlines the hyperparameters used for training.
TABLE VII: Model parameters